OpenStudy (anonymous):

you have 345 ml of 1.5m solution, how much 18m solution do you need to make make the solution 4m.. Please help, i've been stuck for days.

OpenStudy (anonymous):

the m is molar btw

OpenStudy (dtan5457):

Really not sure what the question is asking....or maybe I'm just clueless. What part of chemistry is this?

OpenStudy (anonymous):

This part is solutions.

OpenStudy (anonymous):

I have an idea of what to do and my teacher says I am on the right track but I am missing something

OpenStudy (dtan5457):

What have you thought of so far?

OpenStudy (anonymous):

So my initial thought was to find moles of the 1.5 M solution. so 345ml X 1.5M= 517.5. Then I thought to find the moles you would need to make this solution 4M, which would be 1380 moles (because 1380 moles/345ml=4M). So I figured I could basically subtract the two moles amounts and divide by the 18M. Making it 862.5/18M=479ml

OpenStudy (anonymous):

47.9ml*

OpenStudy (matt101):

Check your units! Molarity is measured in mol/L, but your volumes are still in mL! If you have a 1.5 M solution, 1 L of that solution contains 1.5 mol of solute. From the question you actually have about ⅓ of a L, but you ended up with over 500 moles! To calculate the volume of 18 M solution needed, you can use a modified version of M1V1=M2V2. Remember, each side of this equation gives moles (concentration x volume), but the equation as I wrote it assumes the number of moles is staying the same from situation 1 to situation 2 (i.e. a simple dilution). We can't use this equation directly because our number of moles is changing (the 18M solution is adding BOTH volume and moles into the mix). Here's how to fix it: We know our final solution is going to be 4 M, so the number of moles in it will be 4 x (0.345 + V), where V is our unknown volume of 18 M solution (remember, it's being added to the 0.345 L that we started with). The number of moles in our final 4 M solution will be the sum of the number of moles contributed from our 0.345 L of 1.5 M solution and whatever volume is added from the 18 M solution. In other words, (1.5 x 0.345) + 18V. Putting this all together: \[(1.5 \times 0.345) + 18V = 4 \times (0.345+V)\]\[0.5175 + 18V = 1.38+4V\]\[14V=0.8625\]\[v=0.0616\] So the volume of 18M solution that needs to be added is 0.0616 L, or 61.6 mL. Adding this to your starting volume of 345 mL gives a total volume of 406.6 mL! If you have any questions please let me know!

OpenStudy (anonymous):

Thank you soo much!! @matt101